Circular No. 6785

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 971208 V. Connaughton, National Academy of Sciences/National Research

Council and Marshall Space Flight Center (MSFC); and R. M. Kippen and R. Preece, University Alabama at Huntsville and MSFC, for the BATSE team; and K. Hurley, University of California at Berkeley, for the Ulysses team, write that an unusually long and smooth single-peaked γ-ray burst (GRB) was detected by BATSE on the Compton Gamma-Ray Observatory on Dec. 8.335 UT (trigger 6526) and was detected also by Ulysses. A duration of nearly 800 s makes it longer by an order of magnitude than any other single-peaked GRB observed by BATSE. It is also the longest single episode of emission in a BATSE GRB. Its intensity rose over an interval of 60 s to a maximum flux of 8.97 ± 0.3 × 10^{−7} erg cm^{−2} s^{−1} (between 25 and 1800 keV; integration time 2.048 s) and decayed with a power law of roughly −0.4 over the following 700 s. The total fluence of the event above 25 keV is estimated to be 1.86 ± 0.03 × 10^{−4} erg cm^{−2}. The BATSE location centroid for the GRB is α = 23^{h}45^{m}50^{s}, δ = +77^{o}56′.4 (equinox 2000.0) with a 1σ statistical error of 1^{o}.2. A preliminary joint BATSE/Ulysses InterPlanetary Network annulus is described by a center at α = 23^{h}25^{m}39^{s}, δ = −11^{o}52′.4, with a radius of 87^{o}.906 (half-width 0^{o}.224). A sky map of the event is available at

http://www.batse.msfc.nasa.gov/~kippen/batserbr/brbr obs.html.

SUPERNOVAE S. Deustua, Lawrence Berkeley Laboratory (LBL); R. McMillan,

Apache Point Observatory; H. Newberg, Fermi National Accelerator Laboratory; P. Nugent, C. Pennypacker, and S. Perlmutter, LBL; and M. Strauss, Princeton University, report redshifts (± 0.002) for the host galaxies of four supernovae announced on IAUC 6782; the following redshifts were estimated from emission features in spectra obtained with the 3.5-m ARC Telescope at Sunspot, NM, on the respective dates: SN 1997eb, Nov. 23 UT, 0.085; SN 1997ec, 23, 0.124; SN 1997ed, 24, 0.152; SN 1997ee, 29, 0.166. SN 1997ee is a normal type-Ia supernova near maximum light, and SN 1997ed may be a type-II event.

SUPERNOVAE 1997dn AND 1997dq Unfiltered CCD magnitude estimates (GSC reference stars) from S.

Moretti and S. Tomaselli, Forlí, Italy: SN 1997dn (cf. IAUC 6763), Dec. 8.02 UT, 16.6; SN 1997dq (cf. IAUC 6770), Dec. 8.04, 16.2.

1997 December 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green